OpenStudy (anonymous):

solve the following system x-3y=-15 -4x-3y=0

OpenStudy (shamil98):

x - 3y = -15 -4x -3y = 0 x - 3y = -15 4x + 3y = 0 5x = -15 x = -3 -3 - 3y = -15 -3y = -12 y = 4